# Metric Distortion in Peer Selection

### v1

@@ 2 Preliminaries
We let $\mathbb{N}$ denote the strictly positive integers and, for $n\in\mathbb{N}$, we write $[n]=\{1,\ldots,n\}$ for the first $n$. A linear order $\succ$ on a set $S$ is a complete, transitive, and antisymmetric binary relation on $S$; we denote the set of all linear orders on $[n]$ by $\mathcal{L}(n)$.
@@
We let (${A\choose k}=\{S\subseteq A\mid|S|=k\}$ denote the feasible committees for a given election; i.e., the set of all subsets of $A$ of size $k$.
@@
A distance metric on $A$ is a function $d\colon A\times A\to\mathbb{R}_{+}$ satisfying (i) $d(a,b)=0$ if and only if $a=b$, (ii) $d(a,b)=d(b,a)$ for every $a,b\in A$, and (iii) $d(a,c)\leq d(a,b)+d(b,c)$ for every $a,b,c\in A$. In this paper, we focus on the line metric: We associate each agent $a\in A$ with a position $x_{a}\in(-\infty,\infty)$, and the metric $d$ is defined by $d(a,b)=|x_{a}-x_{b}|$ for every $a,b\in A$. A metric $d$ is said to be consistent with a ranking profile $\succ\ \in\mathcal{L}^{n}(n)$, denoted as $d\ \rhd\succ$, if for every triple of agents $a,b,c\in A$, the condition $d(a,b)<d(a,c)$ holds whenever $b\succ_{a}c$ [^1] Note that this definition allows for agent-dependent tie-breaking; i.e., when $d(a,b)=d(a,c)$ agent $a$ can rank either $b\succ_{a}c$ or $c\succ_{a}b$, independently of other agents. This assumption makes the problem in principle harder, so that our upper bounds on the distortion remain valid if a common tie-breking rule is employed, and it allows us to construct simpler examples for lower bounds. It is not hard to see that the same lower bounds can be obtained without the assumption: Whenever a metric has ties, distances can be perturbed by a small enough constant $\varepsilon$ so that there are no longer ties and the distortion does not improve. Since $d$ is fully defined by the position vector $x\in(-\infty,\infty)^{A}$, we often refer directly to this vector being consistent with a ranking profile $\succ\ \in\mathcal{L}^{n}(n)$ and denote it by $x\ \rhd\succ$. Likewise, we often exchange $d$ by $x$ in the definitions that follow. Finally, for a fixed election $\mathcal{E}=(A,k,\succ)$, consistent vector of locations $x\in(-\infty,\infty)^{n}$, and interval $I=(y,z)$ with $y<z$, we let $A(I)=\{a\in A\mid x_{a}\in I\}$ denote the agents with locations in $I$. When $I$ is a single point $\bar{x}$, we write $A(\bar{x})$ for the agents located at this point.
@@ 2.1 Computing the Order From an Election
An essential property in line metric settings is the ability to determine the order of agents based on their preferences. This result has been established in prior work. Specifically, Elkind and Faliszewski (2014) and Babashah et al. (2024) demonstrate that if the preference lists of voters are pairwise distinct, it is possible to uniquely determine their ordering on the line, along with the ordering of non-Pareto-dominated alternatives. While their setting differentiates between voters and alternatives, this result naturally extends to our context, where agents serve as both voters and candidates. In fact, this follows from a simpler fact in this context: For any three agents, their relative order on the line can be reconstructed from their preference rankings. We state this result as a lemma, which serves as a foundation for many results in this paper as it guarantees that the order of agents in any election can be uniquely identified.
Lemma 1 (Elkind and Faliszewski (2014), Babashah et al. (2024)).
For every election $\mathcal{E}=([n],k,\succ)$, we can compute a permutation $\pi\colon[n]\to[n]$ of the agents such that, for any consistent position vector $x\in(-\infty,\infty)^{n}$ with $x\ \rhd\succ$, we have either $x_{\pi(1)}\leq x_{\pi(2)}\leq\dots x_{\pi(n-1)}\leq x_{\pi(n)}$ or $x_{\pi(n)}\leq x_{\pi(n-1)}\leq\dots\leq x_{\pi(2)}\leq x_{\pi(1)}$.
For simplicity, whenever we fix an election throughout the paper we will assume w.l.o.g. that the agents are already ordered, i.e., that the permutation $\pi$ stated in the lemma is the identity. Hence, we denote the ordered agents by $1,\ldots,n$ and informally refer to this order as from left to right.
@@ 3 Utilitarian Social Cost
[…]
For larger committee sizes ($k>1$), it becomes necessary to define a way to aggregate voters’ distances to the selected agents. In this section, we study two aggregation rules: one that considers the sum of all distances to selected agents in Section 3.1, and one that considers the distance to the $q$th closest agent in Section 3.2.
@@ 3.1 Utilitarian Additive Cost
[…]
For selecting a committee of size $k\geq 2$, we consider the following voting rule.
@@
Compute the order of the agents $1,\ldots,n$ and return $S=\big{\{}\big{\lfloor}\frac{n-k}{2}\big{\rfloor}+1,\dots,\big{\lfloor}\frac{n}{2}\big{\rfloor}+1,\dots,\big{\lfloor}\frac{n+k}{2}\big{\rfloor}\big{\}}$.
Not that the rule selects $k$ agents, leaving $\big{\lfloor}\frac{n-k}{2}\big{\rfloor}$ unselected agents on the left extreme and $n-\big{\lfloor}\frac{n+k}{2}\big{\rfloor}$ unselected agents on the right extremes. These values are equal if $n-k$ is even; the latter is one unit larger if $n-k$ is odd. On an intuitive level, the rule can be understood as constructed by going through the rank list of the median(s) agent(s), selecting agents in the order reported by them but alternating between those to their left and to their right. This ensures a balanced representation of agents on both sides.
Regarding the selection of larger committees, an important ingredient for our results is that an optimal committee selecting consecutive agents always exists. We state this in the following lemma.
@@
Let $\mathcal{E}=(A,k,\succ)$ with $A=[n]$ and $d$ be as in the statement, and let also $x\ \rhd\succ$ be a consistent position vector defining $d$. The result is trivial if $k=1$, so we assume that $k\geq 2$ in what follows. We first observe that, by Lemma $2$ in Babashah et al. (2024), $\text{SC}(a,A;d)\leq\text{SC}(b,A;d)$ holds for $a,b\in A$ are such that either (1) $a,b\geq\frac{n+1}{2}$ and $a-\frac{n+1}{2}\leq b-\frac{n+1}{2}$, or (2) $a,b\leq\frac{n+1}{2}$ and $\frac{n+1}{2}-a\leq\frac{n+1}{2}-b$. In simple words, if two agents lie on the same side of the median agent(s), the agent closer to them has a lower cost. Thus, there exist $S^{*}\in{A\choose k}$ that minimizes the social cost such that $\{m_{1},m_{2}\}\subseteq S^{*}$, where $m_{1}=\big{\lfloor}\frac{n+1}{2}\big{\rfloor}$ and $m_{2}=\big{\lceil}\frac{n+1}{2}\big{\rceil}$ denote the median agent(s) (note that $m_{1}=m_{2}$ if $n$ is odd). Now, suppose that $S^{*}$ is not consecutive. Since $m_{1},m_{2}\in S^{*}$, there exists an agent $a\notin S^{*}$ and $b\in S^{*}$ such that either (1) $a,b\geq\frac{n+1}{2}$ and $a-\frac{n+1}{2}\leq b-\frac{n+1}{2}$, or (2) $a,b\leq\frac{n+1}{2}$ and $\frac{n+1}{2}-a\leq\frac{n+1}{2}-b$. But then, using the result by Babashah et al. again, we obtain that $\text{SC}((S^{*}\setminus\{b\})\cup\{a\},A;d)\leq\text{SC}(S^{*},A;d)$; i.e., we can exchange $b$ by $a$ and the social cost of the committee does not increase. By repeating this procedure, we reach a committee with consecutive agents and minimum social cost, as claimed in the statement. ∎
@@
The distortion of Median Alternation is at most $\frac{2}{k}\Big{(}n-\sqrt{2\big{\lfloor}\frac{n-k}{2}\big{\rfloor}n}\Big{)}$ for utilitarian additive social cost.
The distortion stated in the theorem ranges between $1$ and $2$, except for the case where $k=n-1$ and $k$ is odd, in which it is equal to $\frac{2n}{n-1}$, making it marginally greater than $2$. The bound is equal to $\frac{2}{k}\big{(}n-\sqrt{(n-k)n}\big{)}$ if $n-k$ is even and to $\frac{2}{k}\big{(}n-\sqrt{(n-k-1)n}\big{)}$ if $n-k$ is odd, so that it is better for even values than for neighboring odd values, with more prominent differences for small $k$. Besides these parity differences, the bound takes values closer to $1$ when $k$ is small and closer to $2$ as $k$ approaches $n$. Figure 1 illustrates the bound for $n=100$ and $k$ between $2$ and $n-1$.
[…]
@@
Let $\mathcal{E}=(A,k,\succ)$, $S$, $x$, $\bar{x}$, $\delta$, $x^{-}$, $x^{+}$, $\succ^{-}$, and $\succ^{+}$ be as in the statement. We denote by $d$, $d^{-}$, and $d^{+}$ the distance metrics associated to $x$, $x^{-}$, and $x^{+}$, respectively.
@@
Combining eqs. 1, 2 and 3, we obtain from the definition of utilitarian social cost that
@@
[…]
@@
+$\delta\,|A(\bar{x})|\,\big{(}|S^{\prime}\cap A((-\infty,\bar{x}))|-|S^{\prime}\cap A((\bar{x},+\infty))|\big{)}.$ + ∞ ) ) | ) .
@@
Since $w>-z$, the second inequality is equivalent to
@@
and we update $x^{\prime}_{b}\leftarrow x^{\prime}_{b}\pm\delta^{*}$ for every $b\in A(\bar{x})$ and the sign that ensures not increasing the distortion $\mathrm{dist}(S,A;x^{\prime})$ of $S$. Note that the definition of $\delta^{*}$ ensures both the existence of this sign, due to Lemma 3, and the fact that the number of different positions |$|\{y\in(-\infty,\infty)\mid\exists a\in[n]:x^{\prime}_{a}=y\}|$ is reduced in each step. Thus, the procedure terminates with a vector $x^{\prime}\in(-\infty,\infty)$ such that (1) $x^{\prime}_{a}\in\{x_{1},x_{n}\}$ for every $a\in A$, and (2) the distortion of $S$ under the resulting metric has not decreased. Note that, since the order of the agents has not been changed besides ties, we have either $S^{\prime}=S$ if the committee selected by Median Alternation has not changed or $S^{\prime}\neq S$ but $\text{SC}(S^{\prime},A,x^{\prime})=\text{SC}(S,A;x^{\prime})$ if the committee has changed due to a different tie-breaking. ∎
@@
denote the sets of agents located at positions $x_{1}$ and $x_{n}$ under the position vector $x^{\prime}$, respectively. We let $S_{1}=S^{\prime}\cap A_{1}$ and $S_{2}=S^{\prime}\cap A_{2}$ denote the agents selected by Median Alternation on $\mathcal{E}^{\prime}$ from agents in $A_{1}$ and $A_{2}$, respectively. Then, the social cost of $S^{\prime}$ is given by
@@
$\mathrm{dist}(f)\leq\frac{\text{SC}(S,A;d^{\prime})}{\text{SC}(S^{*},A;d^{\prime})}$ $=\frac{|A_{1}|\cdot|S_{n}|\cdot d^{\prime}(x_{1},x_{n})+|A_{n}|\cdot|S_{1}|\cdot d^{\prime}(x_{1},x_{n})}{|A_{1}|\cdot k\cdot d^{\prime}(x_{1},x_{n})}$
@@
If $|S_{n}|=k$, we obtain $\mathrm{dist}(f)=1$. In what follows, we thus assume $S_{1}\neq\emptyset$. From the definition of the Median Alternation voting rule, we know that $|A_{n}|-|S_{n}|=|A_{1}|-|S_{1}|$ if $n-k$ is even, and either $|A_{n}|-|S_{n}|=|A_{1}|-|S_{1}|+1$ or $|A_{n}|-|S_{n}|=|A_{1}|-|S_{1}|-1$ if $n-k$ is odd. Since the distortion increases in $|S_{1}|$ for fixed $n$ and $k$ due to the assumption that $|A_{n}|\geq|A_{1}|$, the worst case is $|A_{n}|-|S_{n}|=|A_{1}|-|S_{1}|+1$ when $n-k$ is odd, so we restrict to it in what follows. For ease of notation, we define a value $\chi\in\{0,1\}$, such that $\chi=0$ if $n-k$ is even and $\chi=1$ if $n-k$ is odd, so that we can express the previous equations simply as
$|A_{n}|-|S_{n}|=|A_{1}|-|S_{1}|+\chi.$
@@
$|A_{1}|=n-|A_{n}|,\qquad|S_{1}|=\frac{n+k+\chi}{2}-|A_{n}|,\qquad|S_{n}|=|A_{n}|-\frac{n-k-\chi}{2}.$
Replacing in eq. 7, we obtain
$\mathrm{dist}(f)\leq\frac{(n-|A_{n}|)\big{(}|A_{n}|-\frac{n-k-\chi}{2}\big{)}+|A_{n}|\big{(}\frac{n+k+\chi}{2}-|A_{n}|\big{)}}{(n-|A_{n}|)k}$ $=\frac{1}{k}\bigg{(}2|A_{n}|-\frac{(n-k-\chi)n}{2(n-|A_{n}|)}\bigg{)}$
@@
$h^{\prime}(y)=\frac{1}{k}\bigg{(}2-\frac{(n-k-\chi)n}{2(n-y)^{2}}\bigg{)},\qquad h^{\prime\prime}(y)=-\frac{(n-k-\chi)n}{k(n-y)^{3}}.$
@@
$h^{\prime}(y^{*})=0\Longleftrightarrow y^{*}=n-\frac{1}{2}\sqrt{(n-k-\chi)n}.$
Combining this fact with eq. 8, we conclude that
$\mathrm{dist}(f)\leq h(y^{*})$ $=\frac{1}{k}\bigg{(}2\bigg{(}n-\frac{1}{2}\sqrt{(n-k-\chi)n}\bigg{)}-\frac{(n-k-\chi)n}{2\cdot\frac{1}{2}\sqrt{(n-k-\chi)n}}\bigg{)}$
$=\frac{2}{k}\Big{(}n-\sqrt{(n-k-\chi)n}\Big{)},$

### v2

@@ 2 Preliminaries
We let $\mathbb{N}$ denote the strictly positive integers and, for $n\in\mathbb{N}$, we write $n]=\{1,\ldots,n\}$ for the first $n$. A linear order $\succ$ on a set $S$ is a complete, transitive, and antisymmetric binary relation on $S$; we denote the set of all linear orders on $[n]$ by $\mathcal{L}(n)$.
@@
We let ${A\choose k}=\{S\subseteq A\mid|S|=k\}$ denote the feasible committees for a given election; i.e., the set of all subsets of $A$ of size $k$.
@@
A distance metric on $A$ is a function $d\colon A\times A\to\mathbb{R}_{+}$ satisfying (i) $d(a,a)=0$, (ii) $d(a,b)=d(b,a)$ for every $a,b\in A$, and (iii) $d(a,c)\leq d(a,b)+d(b,c)$ for every $a,b,c\in A$. In this paper, we focus on the line metric: We associate each agent $a\in A$ with a position $x_{a}\in(-\infty,\infty)$, and the metric $d$ is defined by $d(a,b)=|x_{a}-x_{b}|$ for every $a,b\in A$. A metric $d$ is said to be consistent with a ranking profile $\succ\ \in\mathcal{L}^{n}(n)$, denoted as $d\ \rhd\succ$, if for every triple of agents $a,b,c\in A$, the condition $d(a,b)<d(a,c)$ implies $b\succ_{a}c$. [^1] Note that this definition allows for agent-dependent tie-breaking; i.e., when $d(a,b)=d(a,c)$ agent $a$ can rank either $b\succ_{a}c$ or $c\succ_{a}b$, independently of other agents. This assumption makes the problem in principle harder, so that our upper bounds on the distortion remain valid if a common tie-breaking rule is employed, and it allows us to construct simpler examples for lower bounds. It is not hard to see that the same lower bounds can be obtained without the assumption: Whenever a metric has ties, distances can be perturbed by a small enough constant $\varepsilon$ so that there are no longer ties and the distortion does not improve. Since $d$ is fully defined by the position vector $x\in(-\infty,\infty)^{A}$, we often refer directly to this vector being consistent with a ranking profile $\succ\ \in\mathcal{L}^{n}(n)$ and denote it by $x\ \rhd\succ$. Likewise, we often exchange $d$ by $x$ in the definitions that follow. Finally, for a fixed election $\mathcal{E}=(A,k,\succ)$, consistent vector of locations $x\in(-\infty,\infty)^{n}$, and interval $I=(y,z)$ with $y<z$, we let $A(I)=\{a\in A\mid x_{a}\in I\}$ denote the agents with locations in $I$. When $I$ is a single point $\bar{x}$, we write $A(\bar{x})$ for the agents located at this point.
@@ 2.1 Computing the Order From an Election
An essential property in line metric settings is the ability to determine the order of agents based on their preferences. This result has been established in prior work. Specifically, Elkind and Faliszewski (2014) and Babashah et al. demonstrate that if voters’ preference lists are pairwise distinct, it is possible to uniquely determine the ordering on the line, along with the ordering of the non-Pareto-dominated alternatives. While their setting differentiates between voters and alternatives, this result naturally extends to our context, where agents serve as both voters and candidates. In our setting, this follows directly from a simpler observation: for any three agents, their relative order on the line can be reconstructed from their preference rankings. We state this result as a lemma, which serves as a foundation for many results in this paper as it guarantees that the order of agents in any election can be uniquely identified.
Lemma 1 (Elkind and Faliszewski (2014), Babashah et al. ).
For every election $\mathcal{E}=([n],k,\succ)$, we can compute a permutation $\pi\colon[n]\to[n]$ of the agents such that, for any consistent position vector $x\in(-\infty,\infty)^{n}$ with $x\ \rhd\succ$, we have either $x_{\pi(1)}\leq x_{\pi(2)}\leq\dots\leq x_{\pi(n)}$ or $x_{\pi(n)}\leq x_{\pi(n-1)}\leq\dots\leq x_{\pi(1)}$.
We briefly highlight some structural features that distinguish the peer selection setting from the general setting. First, unlike the general voting setup with disjoint voters and candidates, the peer selection problem provides us with not only the order of the candidates but also the exact order of the voters. Second, this order can be efficiently reconstructed: we can identify extreme agents by observing who consistently appears at the bottom of others’ rankings. Once such an agent is identified, their ranking determines the full order of agents along the line. Third, the mutual evaluations are more restricted, so we gain more information about relative preferences, such as how many voters prefer one candidate over another, than in the general voting setting. These features inform the design of improved mechanisms in peer selection and present new algorithmic and structural challenges.
For simplicity, whenever we fix an election throughout the paper, we will assume w.l.o.g., that the agents are already ordered, i.e., that the permutation $\pi$ stated in the lemma is the identity. Hence, we denote the ordered agents by $1,\ldots,n$ and informally refer to this order as from left to right.
@@ 3 Utilitarian Social Cost
Using Lemma 1, we know that the order of agents can be fully determined from the preference profile $\mathbf{\succ}$. This allows us to identify the median agent, who minimizes the total distance to all other agents. Hence, when selecting a single agent ($k=1$) under the utilitarian objective, the median agent is optimal. This observation holds regardless of whether the individual cost is defined additively or via the $q$-cost, since both notions coincide when $k=1$.
For larger committee sizes ($k>1$), it becomes necessary to define how a voter’s distances to the selected agents are aggregated. In this section, we study two aggregation rules: one that considers the sum of all distances to selected agents in Section 3.1, and one that considers the distance to the $q$th closest selected agent in Section 3.2.
@@ 3.1 Utilitarian Additive Cost
It is straightforward to verify that the optimal committee can be directly identified from the preference profile when the committee size is both $k=1$ and $k=2$. For $k=1$, the optimal committee consists of the median agent, as discussed above. For $k=2$, the optimal committee consists of the two median agents if $n$ is even, or the median agent together with the agent closest to them, if $n$ is odd. In both cases, the optimal agents can be determined solely from $\mathbf{\succ}$, without knowledge of the underlying metric, yielding a distortion of $1$.
We now introduce our voting rule for selecting a committee of size $k\geq 2$, called the Median Alternation rule. This rule can be viewed as a generalization of the Polar Comparison rule proposed by Babashah et al. for multi-winner elections. In contrast to the Polar Comparison rule, the Median Alternation rule (i) applies to all committee sizes $k$, (ii) is simpler since it does not require case distinctions or pairwise comparisons, and (iii) achieves improved distortion guarantees under peer selection.
@@
Compute the order of the agents $1,\ldots,n$ on the line. If nkn-k is even, return $S=\big\{\tfrac{n-k}{2}+1,\;\tfrac{n-k}{2}+2,\;\dots,\;\tfrac{n+k}{2}\big\}$. If nkn-k odd and $\big\lceil\tfrac{n-k}{2}\big\rceil\succ_{m}\big\lceil\tfrac{n+k}{2}\big\rceil$, return $S=\big\{\big\lceil\tfrac{n-k}{2}\big\rceil,\,\big\lceil\tfrac{n-k}{2}\big\rceil+1,\,\dots,\,\big\lceil\tfrac{n+k}{2}\big\rceil-1\big\}$. Otherwise, return $S=\big\{\big\lceil\tfrac{n-k}{2}\big\rceil+1,\,\big\lceil\tfrac{n-k}{2}\big\rceil+2,\,\dots,\,\big\lceil\tfrac{n+k}{2}\big\rceil\big\}$.
On an intuitive level, the rule can be understood as constructed by going through the rank list of the median(s) agent(s), selecting agents in the order reported by them but alternating between those to their left and to their right. This ensures a balanced representation of agents on both sides.
For larger committees, a key ingredient in our analysis is that there always exists an optimal committee consisting of consecutive agents. We formalize this observation in the following lemma.
@@
Let $\mathcal{E}=(A,k,\succ)$ with $A=[n]$ and $d$ be as in the statement, and let also $x\ \rhd\succ$ be a consistent position vector defining $d$. The result is trivial if $k=1$, so we assume that $k\geq 2$ in what follows. We first observe that, by Lemma $2$ in Babashah et al. , $\text{SC}(a,A;d)\leq\text{SC}(b,A;d)$ holds for $a,b\in A$ are such that either (1) $a,b\geq\frac{n+1}{2}$ and $a-\frac{n+1}{2}\leq b-\frac{n+1}{2}$, or (2) $a,b\leq\frac{n+1}{2}$ and $\frac{n+1}{2}-a\leq\frac{n+1}{2}-b$. In simple words, if two agents lie on the same side of the median agent(s), the agent closer to them has a lower cost. Thus, there exist $S^{*}\in{A\choose k}$ that minimizes the social cost such that $\{m_{1},m_{2}\}\subseteq S^{*}$, where $m_{1}=\big\lfloor\frac{n+1}{2}\big\rfloor$ and $m_{2}=\big\lceil\frac{n+1}{2}\big\rceil$ denote the median agent(s) (note that $m_{1}=m_{2}$ if $n$ is odd). Now, suppose that $S^{*}$ is not consecutive. Since $m_{1},m_{2}\in S^{*}$, there exists an agent $a\notin S^{*}$ and $b\in S^{*}$ such that either (1) $a,b\geq\frac{n+1}{2}$ and $a-\frac{n+1}{2}\leq b-\frac{n+1}{2}$, or (2) $a,b\leq\frac{n+1}{2}$ and $\frac{n+1}{2}-a\leq\frac{n+1}{2}-b$. But then, using the result by Babashah et al. again, we obtain that $\text{SC}((S^{*}\setminus\{b\})\cup\{a\},A;d)\leq\text{SC}(S^{*},A;d)$; i.e., we can exchange $b$ by $a$ and the social cost of the committee does not increase. By repeating this procedure, we reach a committee with consecutive agents and minimum social cost, as claimed in the statement. ∎
@@
For every $n,k\in\mathbb{N}$ with $n\geq k\geq 2$, under the utilitarian additive social cost the distortion of Median Alternation is at most
$\frac{1}{k}\left(2n+\chi-2\sqrt{\,n\bigl(n-k+\chi\bigr)}\right),$
where $\chi=1$ if nkn-k is odd and $\chi=0$ otherwise.
The bound ranges between $1$ and $2$: it is closer to $1$ when $k$ is small relative to $n$, and approaches $2$ as $k$ approaches $n$. Figure 1 illustrates the bound for $n=100$ and $k$ between $2$ and n1n-1.
11kk0distortion1010202030304040505060607070808090901001001.21.21.41.41.61.61.81.822
@@
Let $\mathcal{E}=(A,k,\succ)$, $S$, $x$, $\bar{x}$, $\delta$, $^{-}, x+x^{+}$, $\succ^{-}$, and $\succ^{+}$ be as in the statement. We denote by $d$, $d^{-}$, and $d^{+}$ the distance metrics associated to $x$, $x^{-}$, and $x^{+}$, respectively.
@@
Combining inequalities 1, 2 and 3, we obtain from the definition of utilitarian social cost that
@@
$\phantom{{}={}}-\delta\,|A(\bar{x})|\,\big(|S^{\prime}\cap A((-\infty,\bar{x}))|-|S^{\prime}\cap A((\bar{x},+\infty))|\big).$
@@
$\phantom{{}={}}+\delta\,|A(\bar{x})|\,\big(|S^{\prime}\cap A((-\infty,\bar{x}))|-|S^{\prime}\cap A((\bar{x},+\infty))|\big).$
@@
Since w>zw>-z, the second inequality is equivalent to
@@
and we update $x^{\prime}_{b}\leftarrow x^{\prime}_{b}\pm\delta^{*}$ for every $b\in A(\bar{x})$ and the sign that ensures not increasing the distortion $\mathrm{dist}(S,A;x^{\prime})$ of $S$. Note that the definition of $\delta^{*}$ ensures both the existence of this sign, due to Lemma 3, and the fact that the number of different positions $|\{y\in(-\infty,\infty)\mid\exists a\in[n]:x^{\prime}_{a}=y\}|$ is reduced in each step. Thus, the procedure terminates with a vector $x^{\prime}\in(-\infty,\infty)$ such that (1) $x^{\prime}_{a}\in\{x_{1},x_{n}\}$ for every $a\in A$, and (2) the distortion of $S$ under the resulting metric has not decreased. Note that, since the order of the agents has not been changed besides ties, we have either $S^{\prime}=S$ if the committee selected by Median Alternation has not changed or $S^{\prime}\neq S$ but $\text{SC}(S^{\prime},A,x^{\prime})=\text{SC}(S,A;x^{\prime})$ if the committee has changed due to a different tie-breaking. ∎
@@
denote the sets of agents located at positions $x_{1}$ and $x_{n}$ under the position vector $x^{\prime}$, respectively. We let $S_{1}=S^{\prime}\cap A_{1}$ and $S_{n}=S^{\prime}\cap A_{n}$ denote the agents selected by Median Alternation on $\mathcal{E}^{\prime}$ from agents in $A_{1}$ and $A_{n}$, respectively. Then, the social cost of $S^{\prime}$ is given by
@@
$\mathrm{dist}(f)$ $\leq\frac{\text{SC}(S,A;d^{\prime})}{\text{SC}(S^{*},A;d^{\prime})}$
$=\frac{|A_{1}|\cdot|S_{n}|\cdot d^{\prime}(x_{1},x_{n})+|A_{n}|\cdot|S_{1}|\cdot d^{\prime}(x_{1},x_{n})}{|A_{1}|\cdot k\cdot d^{\prime}(x_{1},x_{n})}$
@@
If $|S_{n}|=k$, we obtain $\mathrm{dist}(f)=1$. In what follows, we thus assume $S_{1}\neq\emptyset$. From the definition of the Median Alternation voting rule, we know that $|A_{n}|-|S_{n}|=|A_{1}|-|S_{1}|$ if nkn-k is even, and $|A_{n}|-|S_{n}|=|A_{1}|-|S_{1}|-1$ if nkn-k is odd.
We can express the previous equations simply as
$|A_{n}|-|S_{n}|=|A_{1}|-|S_{1}|-\chi.$
@@
$|A_{1}|=n-|A_{n}|,\quad|S_{1}|=\frac{n+k-\chi}{2}-|A_{n}|,\quad|S_{n}|=|A_{n}|-\frac{n-k-\chi}{2}.$
Replacing in inequality 7, we obtain
$\mathrm{dist}(f)$ $\leq\frac{(n-|A_{n}|)\big(|A_{n}|-\frac{n-k-\chi}{2}\big)+|A_{n}|\big(\frac{n+k-\chi}{2}-|A_{n}|\big)}{(n-|A_{n}|)k}$
$=\frac{2|A_{n}|}{k}-\frac{\chi|A_{n}|}{k(n-|A_{n}|)}-\frac{n(n-k-\chi)}{2k(n-|A_{n}|)}$
@@
$h^{\prime}(y)=\frac{1}{k}\bigg(2-\frac{(n-k+\chi)n}{2(n-y)^{2}}\bigg),\qquad h^{\prime\prime}(y)=-\frac{(n-k+\chi)n}{k(n-y)^{3}}.$
@@
$h^{\prime}(y^{*})=0\Longleftrightarrow y^{*}=n-\frac{1}{2}\sqrt{(n-k+\chi)n}.$
Combining this fact with inequality 8, we conclude that
$\mathrm{dist}(f)$ $\leq h(y^{*})$
$=\frac{2}{k}\bigg(n-\frac{1}{2}\sqrt{(n-k+\chi)n}\bigg)-\frac{\chi\big(n-\frac{1}{2}\sqrt{(n-k+\chi)n}\big)}{\frac{k}{2}\sqrt{(n-k+\chi)n}}$
$\phantom{{}={}}-\frac{n(n-k-\chi)}{k\sqrt{(n-k+\chi)n}}$
$=\frac{1}{k}\Big(2n+\chi-2\sqrt{(n-k+\chi)n}\Big),$